OpenStudy (anonymous):

A 500 ohm resistor and a 1.2-mH inductor are connected in parallel to a 12-V,40-kHz source. How do I find the phase angle (theta)?

OpenStudy (kropot72):

 \[\theta=\tan^{-1} \frac{I _{L}}{I _{R}}=\tan^{-1} \frac{\frac{V}{\omega L}}{\frac{V}{R}}=\tan^{-1} \frac{R}{\omega L}\]